If a mixed-design experiment contains four levels of the between-groups factor and six levels of the repeated-measures factor, and there are 15 participants in each group, how many degrees of freedom are associated with the error term for the repeated-measures factor?

a) 224

b) 280

c) 210

d) 300
Answer #1

If a mixed-design experiment contains four levels of the between-groups factor and six levels of the repeated-measures factor, and there are 15 participants in each group.

That is, i =4, j =6 and n = 15

The degrees of freedom are associated with the error term for the repeated-measures factor is,

DF = (i - 1) * (j - 1) * (n - 1)

=> DF = (4 - 1) * (6 - 1) * (15 - 1)

=> DF = 3 * 5 * 14

=> DF = 210

Error degrees of freedom = 210